On 1 July 20X8, Sun Company purchased $4,000,000 of Moon Corp. 6.2% bonds, classified as an AC investment. The bonds pay semi-annual interest each 30 June and 31 December. The market interest rate was 6% on the date of purchase. The bonds mature on 30 June 20X13.

Required:
1. Calculate the price paid by Sun Company.
2. Construct a table that shows interest revenue reported by Sun, and the carrying value of the investment, for each interest period for four interest periods. Use the effective-interest method.

Period

Cash Payment

Interest Revenue

Amortization

Bond Carrying Value

1

2

3

....


3. Prepare the entries for the first two interest periods based on your calculations in requirement 2.

4. How much interest revenue would be reported for the year ended 31 December 20X9?

5. What would the balance of the investment account for the year ended 31 December 20X9?
Answer #1

Given details in the question are,

Bond maturity = 5years

Bond Face value = $4000000

Coupon rate of interest = 6.2% per annum and interest payable semi annually.

Total period = 5 years *2= 10 terms and

Hence, coupon rate per teem = 3.1% per term

Market rate of interest = 6% per annum and 3% per term

1.Computing the bond value:

Term Particulars amount PVF@3% PV
1-10 Coupon $4000000*3.1%=124000 8.5302 1057744.80
10 Maturity amount $4000000 0.7440 2976000
Present value of Bond 4033744.80

As the Coupon rate is more than the market rate, the bond is selling at premium

Premium = 4033744.80-4000000= $33744.80

2. Amortization table using Effective intereat method for 4 interest periods:

Date(A) Interest payment(B) Interest Expense(C=G*3%) Amortisation(D=B-C) Bond premium(E) Bonds payable(F) Carrying amount(G=E+F)
1 july 20X8 33744.80 4000000 4033744.80
31 Dec 20X8 124000 121012.34 2987.66 30757.14 4000000 4030757.14
30 june 20X9 124000 120922.71 3077.29 27679.85 4000000 4027679.85
31 Dec 20X9 124000 120830.39 3169.60 24510.24 4000000 4024510.24
30 june 20X0 124000 120735.30 3264.70 21245.54 4000000 4021245.54

3. Journal Entris for the first two interes periods:

1 july 20X8 Cash Dr 4033744.80

To Bonds payable 4000000

To Premium on Bonds payable 33744.80

(At the time of issue)

31 Dec 20X8 Interest expense Dr 121012.34

Premium on Bonds payable Dr 2987.66

To Cash 124000

(At the time of Interest payment)

30 june 20X9   Interest expense Dr 120922.71

Premium on Bonds payable Dr 3077.29

To Cash 124000

(At the time of Interest payment)

4.The Interest revenue to be reported for the year ended 31 december 20X9 = (241753.1) which is the total of the intersst paid during the year as mentioned in the table above.

5. Balance of the investment account for the year ended 31 December 20X9 = $4024510.24 as per the table above.